1. What is emergent behavior? Give an example of emergent behavior.

2. Imagine a cart of mass m is traveling at a constant speed v. A mass of 3 m is dropped vertically onto the cart. What is the speed of the cart after contact? ''

3. A ball of mass m is thrown at a wall and bounces off at ½ its initial velocity (in the opposite direction). What is the change in momentum?

4. Consider a cart on a track traveling with a constant (but slightly decreasing velocity). Explain how impulse might be used to explain the change in speed/velocity/momentum of the cart.
Answer #1

1) Emergent behavior is the one which does not depend on individual but a group with a strong bond and relationship. The perfect example for this is a large group of ants getting together to build complex structures and colonies

----------------

2) This is an inelastic collision

the 3m mass was at rest initially

so,

vf = mv / m + 3m

vf = mv / 4m

vf = v/4

---------------------------

3) change in momentum = mass ( final velocity - initial velocity)

I am assuming the object was initially moving in positive direction

then

change in momentum = m ( - 1/2v - v)

change in momentum = -1.5mv

---------------------------------

4) impulse = change in momentum